# Low-Confidence Gold: Refining Low-Confidence Samples for Efficient Instruction Tuning

### v4

@@ Abstract
The effectiveness of instruction fine-tuning for Large Language Models is fundamentally constrained by the quality and efficiency of training datasets. This work introduces Low-Confidence Gold (LCG), a novel filtering framework that employs centroid-based clustering and confidence-guided selection for identifying valuable instruction pairs. Through a semi-supervised approach using a lightweight classifier trained on representative samples, LCG curates high-quality subsets while preserving data diversity. Experimental evaluation demonstrates that models fine-tuned on LCG-filtered subsets of 6K samples achieve superior performance compared to existing methods, with substantial improvements on MT-bench and consistent gains across comprehensive evaluation metrics. The framework’s efficacy while maintaining model performance establishes a promising direction for efficient instruction tuning.
@@
Hongyi Cai1, Jie Li2, Mohammad Mahdinur Rahman1, Wenzhen Dong3 1Universiti Malaya, 2University of Science and Technology Beijing, 3The Chinese University of Hong Kong Correspondence: xcloudfance@gmail.com
@@ 1 Introduction
Large Language Models (LLMs) have been trained to follow instructions by specific supervised response data after pre-training stage. Many instruction finetuning (IFT) Taori et al. (2023) datasets emerge to realize various downstream tasks, for example: mathematic calculation, sentence analysis, haiku writing and etc, aiming to strengthen the ability of LLMs in instruction following. To save vast human costs for data annotation, most of studies introduce other teacher LLMs (e.g. text-davinci-003 Brown et al. (2020)) to align the best instructions with corresponding responses.
However, IFT datasets (e.g. Alpaca_52k Taori et al. (2023), magpie Xu et al. (2024)) suffer from misleading content and poor quality, resulting in the bottleneck of post-training performance, even though teacher models replenish the missing parts of context and instruction pairs. This highlights the need for effective data filtering methods that identify high-quality instruction subsets while reducing fine-tuning time and computational costs.
@@
Alpagasus Chen et al. (2024) proposed a model-based approach that introduces proprietary LLMs to score data quality in multiple facets, replacing human annotation by taking advantage of the automated pipeline. However, this leads to datasets that are likely biased by the preference for redundant and limited responses Panickssery et al. (2024), which potentially deteriorates the diversity of the original data. Ge et al., 2024 emphasizes the necessity of diversity and therefore proposed clustering and ranking to select subsets of data. Further, Superfiltering Li et al. (2024) gains more insights in small open-source LLM that scores the instruction following ability of Alpaca_52k. Although the instruction score provides an efficient and simple criterion for data selection, it does not consistently correlate with both the quality and diversity of data. Consequently, improvements in performance may not always be guaranteed.
@@ 3.1 Motivation
Instruction filtering demands a dual-focus mechanism that intrinsically balances data quality and diversity. Traditional supervised methods face inherent scalability limitations as manual annotation becomes prohibitively expensive for large-scale instruction datasets Liu et al. (2022); Longpre et al. (2023); Liu et al. (2023). Meanwhile, it is difficult to identify suitable and challenging data for LLMs training without introducing proprietary LLMs or labors. Our semi-supervised framework addresses these limitations through pseudo-label refinement and early-stopped confidence detection, creating dynamic selection boundaries aligned with language model learning dynamics.
@@
Table 1: Performance comparison on standard benchmarks. Results in bold indicate best performance within each group, while underlined values represent second-best performance within each group. The table is divided into two groups, each with its base model and various fine-tuning methods.
3.2 Centroid Coreset Selection for Pseudo-labels
In the initial step of our approach, we select a coreset from the whole corpus to identify pseudo-labels by the K-means algorithm, which effectively determine each semantic clusters. Given a dataset of instruction pairs $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, we first encode each instruction $\mathbf{x}_{i}$ into a dense vector representation using MiniLM Wang et al. (2022):
where $\gamma$ is the 90th percentile distance within each cluster.
$\cdot\log p_{\theta}(y=j|\mathbf{x}_{i})$
@@ 4 Experiments
In this section, we utilize LCG to filter Alpaca_52k dataset into 6k and evaluate the subset by fine-tuning in 2 open-source LLMs: 1) Mistral-7b-v0.3 Jiang et al. (2023) and 2) LLaMa3-8b Dubey et al. (2024). Additionally, we adopt two different classifiers as our semi-supervised training model to examine difficult samples. Both results demonstrate the effectiveness of our method, as shown in Tab. 1.
Settings. Two classifiers are Multinomial Naive Bayes Pedregosa et al. (2011) and DistilBERT Sanh et al. (2019) respectively. We set the training learning rate as 1e-5 and also 3 epochs to train as mentioned before. After training is finished, we select curated datasets by confidences < 0.7 to fine-tune open-source LLMs by LoRA Hu et al. (2022) with a learning rate of 2e-5 and 3 epochs.
Our proposed Low-confidence Gold (LCG) method consistently outperforms existing instruction data filtering approaches across multiple base models and evaluation benchmarks. When applied to Mistral-7b, LCG with MultinomialNB achieves the highest MT-bench score of 5.086, surpassing the previous best (LIMA-6k Zhao et al. (2024)) by 14.5%. Similarly, LCG with DistilBERT demonstrates superior performance on LLaMA3-8b, improving the MT-bench score by 11.5% over LIMA-6k. Notably, our method maintains strong performance even with only 1k examples, highlighting its effectiveness in identifying high-quality instruction data. The consistent improvements across diverse metrics (Hellaswag Zellers et al. (2019), MMLU Hendrycks et al. (2021), GSM8k Cobbe et al. (2021), and ARC Clark et al. (2018)) further validate the robustness of our approach.

### v6

@@ Abstract
The effectiveness of instruction fine-tuning for Large Language Models is fundamentally constrained by the quality and efficiency of training datasets. This work introduces Low-Confidence Gold (LCG), a novel filtering framework that employs centroid-based clustering and confidence-guided selection for identifying valuable instruction pairs. Through a semi-supervised approach using a lightweight classifier trained on representative samples, LCG curates high-quality subsets while preserving data diversity. Experimental evaluation demonstrates that models fine-tuned on LCG-filtered subsets of 6K samples achieve superior performance compared to existing methods, with substantial improvements on MT-bench and consistent gains across comprehensive evaluation metrics. The framework’s efficacy while maintaining model performance establishes a promising result for efficient instruction tuning.
@@
Hongyi Cai1,*   Jie Li2,*   Mohammad Mahdinur Rahman1   Wenzhen Dong3 1Universiti Malaya  2University of Science and Technology Beijing  3The Chinese University of Hong Kong *Corresponding: xcloudfance@gmail.com
@@ 1 Introduction
Large Language Models (LLMs) have been trained to follow instructions by specific supervised response data after pre-training stage. Many instruction finetuning (IFT) alpaca datasets emerge to realize various downstream tasks, for example: mathematic calculation, sentence analysis, haiku writing and etc, aiming to strengthen the ability of LLMs in instruction following. To save vast human costs for data annotation, most of studies introduce other teacher LLMs (e.g. text-davinci-003 gpt) to align the best instructions with corresponding responses.
However, IFT datasets (e.g. Alpaca_52k alpaca, magpie xu2024magpie) suffer from misleading content and poor quality, resulting in the bottleneck of post-training performance, even though teacher models replenish the missing parts of context and instruction pairs. This highlights the need for effective data filtering methods that identify high-quality instruction subsets while reducing fine-tuning time and computational costs.
@@
Alpagasus chen2023alpagasus proposed a model-based approach that introduces proprietary LLMs to score data quality in multiple facets, replacing human annotation by taking advantage of the automated pipeline. However, this leads to datasets that are likely biased by the preference for redundant and limited responses panickssery2024llm; mergeit, which potentially deteriorates the diversity of the original data. ge2024clustering emphasizes the necessity of diversity and therefore proposed clustering and ranking to select subsets of data. Further, Superfiltering li2024superfiltering gains more insights in small open-source LLM that scores the instruction following ability of Alpaca_52k. Although the instruction score provides an efficient and simple criterion for data selection, it does not consistently correlate with both the quality and diversity of data. Consequently, improvements in performance may not always be guaranteed.
@@ 3.1 Motivation
Instruction filtering demands a dual-focus mechanism that intrinsically balances data quality and diversity. Traditional supervised methods face inherent scalability limitations as manual annotation becomes prohibitively expensive for large-scale instruction datasets liu-etal-2022-makes; longpre2023flan; deita. Meanwhile, it is difficult to identify suitable and challenging data for LLMs training without introducing proprietary LLMs or labors. Our semi-supervised framework addresses these limitations through pseudo-label refinement and early-stopped confidence detection, creating dynamic selection boundaries aligned with language model learning dynamics.
@@
3.2 Centroid Coreset Selection for Pseudo-labels
In the initial step of our approach, we select a coreset from the whole corpus to identify pseudo-labels by the K-means algorithm, which effectively determine each semantic clusters. Given a dataset of instruction pairs $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, we first encode each instruction $\mathbf{x}_{i}$ into a dense vector representation using MiniLM wang2022minilmv2:
where $\gamma$ is the 90th percentile distance within each cluster.
$\cdot\log p_{\theta}(y=j|\mathbf{x}_{i})$
@@ 3.3 Low-Confidence Gold: Calibrating with Low-confidence samples to select data
Table 1: Performance comparison on standard benchmarks. Results in bold indicate best performance within each group, while underlined values represent second-best performance within each group. The table is divided into two groups, each with its base model and various fine-tuning methods. We add the complete results for LCG-MultinomialNB-6k on LLaMa3-8b.
@@ 4 Experiments
4.1 Experimental Setup
We utilize LCG to filter the Alpaca_52k dataset. Two classifiers are used for the selection process: Multinomial Naive Bayes (MNB) scikit-learn and DistilBERT sanh2019distilbert. The classifiers are trained for 3 epochs with a learning rate of 1e-5. We then select curated datasets by a confidence threshold of <0.7<0.7. The resulting subsets are used to fine-tune two open-source LLMs: Mistral-7b-v0.3 jiang2023mistral and LLaMa3-8b dubey2024llama, using LoRA hu2021lora with a learning rate of 2e-5 for 3 epochs.
Models (Mistral-7B) ARC GSM8k HellaSwag MMLU Avg
WizardLM-50k (Full) 51.54 38.59 62.14 59.36 52.91
WizardLM-Longest 51.08 38.99 62.10 58.99 52.79
WizardLM-Perplexity 51.19 39.77 62.21 59.01 53.05
WizardLM-K-means 51.96 39.95 62.11 59.38 53.35
WizardLM-SuperFiltering 51.54 39.65 62.21 59.22 53.16
LCG-DistilBERT-1k (Ours) 52.31 40.00 62.39 60.01 53.68
LCG-MNB-6k (Ours) 51.67 38.97 62.28 60.75 53.42
Table 2: Cross-dataset validation on WizardLM with Mistral-7B. Our LCG method demonstrates superior performance against baselines.
4.2 Main Results
As presented in Table 1, our proposed LCG method consistently outperforms existing instruction data filtering approaches. When applied to Mistral-7b, LCG with MultinomialNB achieves the highest MT-bench score of 5.086, surpassing the previous best (LIMA-6k LIMA) by 14.5%. Similarly, LCG with DistilBERT demonstrates superior performance on LLaMA3-8b, improving the MT-bench score by 11.5% over LIMA-6k. Notably, our method maintains strong performance even with only 1k examples, highlighting its effectiveness. The consistent improvements across diverse metrics (Hellaswag hellaswag, MMLU mmlu, GSM8k gsm8k, and ARC arc) further validate the robustness of our approach.
Our LCG-tuned models show particularly strong performance on MT-Bench. We attribute this success to the fact that low-confidence samples often represent nuanced dialogue scenarios, edge cases, and diverse response patterns. Training on these samples equips the model with superior reasoning capabilities, essential for the sophisticated multi-turn conversations in MT-Bench.
5 Analysis and Discussion
5.1 Cross-Dataset Generalization
To validate the robustness of our approach, we applied LCG to the WizardLM dataset, which is larger and more complex than Alpaca. As shown in Table 2, our method continues to outperform other selection strategies, confirming that LCG is not overfitted to a single dataset but offers a generally applicable mechanism for curating high-quality instruction data.
5.2 Ablation Studies
We conducted ablation studies on key hyperparameters to validate our methodological choices.
Impact of Confidence Threshold $\tau$
The confidence threshold $\tau$ directly controls the size of the filtered subset. As shown in our main results (Table 1), relaxing $\tau$ to expand the dataset from 1K to 6K samples consistently improves performance on reasoning tasks like GSM8k. However, further experiments with a looser threshold (yielding 9K samples) resulted in significant performance degradation, indicating that our moderate threshold effectively balances quality and quantity by filtering out noise.
5.3 Synergy with RLHF
To explore modularity, we designed a two-stage pipeline combining LCG with Reinforcement Learning from Human Feedback (RLHF). First, LCG selects 10K candidates. Then, an RLHF preference model refines this set to the top 6K. As shown in Table 3, this combined approach significantly improves performance, especially on reasoning tasks, demonstrating that LCG serves as an effective initial filter for more advanced refinement techniques.
Method GSM8k HellaSwag MMLU Avg
RLHF-Only-6k 39.85 61.45 60.12 53.84
LCG+RLHF-6k 42.15 62.35 60.45 54.66
Table 3: Performance of combining LCG with RLHF. The synergistic approach yields superior results, especially on reasoning.